# Languages of Boundedly-Ambiguous Vector Addition Systems with States

### v1

@@ 3.2 Semilinear image
Now, we develop the second tool for showing that language is not recognized by a $k$-ambiguous VASS. Before formulating the tool we have to provide a few definitions. For any language $L\subseteq\{a,b\}^{*}$ such that for each $w\in L$ we have #$\#_{b}(w)=l$ for some $l\in\mathbb{N}$
@@
  • $L$ is recognized by $k$-ambiguous VASS $V$.
  • There exists $n\in\mathbb{N}$ such that for each $w\in L$ we have #$\#_{b}(w)=n$.
@@
{toappendix}
Proof 3.24 (Proof of Lemma 3.23).
Notice, that $L$ satisfies conditions of Lemma 3.4. Therefore we can apply Lemma 3.4 to $L$ and $n+1$ and decompose each accepting run in $V$ in the way presented in Lemma 3.4. Let us fix some $\alpha_{1},\ldots,\alpha_{n+1}$, $\alpha_{1}^{\prime},\ldots,\alpha_{n+1}^{\prime}$ and $\beta_{1},\ldots,\beta_{n+1}$, initial configuration $q(c)$ and accepting configuration $p(f)$. We call a run $\rho$ accepting with respect to $q(c)$ and $p(f)$ if $\rho$ is an accepting run of the VASS starting in $q(c)$ and ending in configuration $p(c^{\prime})$ such that $c^{\prime}\geq f$. Let $K$ be the following language:
@@
$,|\beta_{n}|a_{n}+|\alpha_{n}|+|\alpha_{n}^{\prime}|-1,|\beta_{k+1}|a_{n+1}+|\alpha_{n+1}|+|\alpha_{n+1}^{\prime}|)\mid$
$\text{such that $\alpha_{1}\beta_{1}^{a_{1}}\alpha_{1}^{\prime}\alpha_{2}\beta_{2}^{a_{2}}\alpha_{2}^{\prime}\ldots\alpha_{n+1}\beta_{n+1}^{a_{k+1}}\alpha_{n+1}^{\prime}$ is an accepting run with respect to $q(c)$ and $p(f)$}\}$
@@
The proof of Lemma 3.23 is based on Lemma 3.4 and the fact that set of solution of a system of linear Diophantine inequalities is semilinear. It is delegated to the Appendix.
We can extend definition of $im(L)$ to different letters by setting:
$im_{c_{1},c_{2}}(L)=\{(a_{1},a_{2},\ldots,a_{l+1})\mid a_{1},a_{2},\ldots,a_{l+1}\in\mathbb{N},c_{1}^{a_{1}}c_{2}c_{1}^{a_{2}}c_{2}\ldots c_{2}c_{1}^{a_{l+1}}\in L\}$
Then, as shown in Lemma 4.1, $k$-ambiguous are closed under intersection with regular languages. Hence, the following corollary holds:
@@
Let $L\subseteq\Sigma^{*}$ be a language recognized by a $k$-ambiguous VASS over. Then for any $a,b\in\Sigma$ and $n\in N$ we have that $im_{a,b}(L\cap L((a^{*}b)^{n}a^{*}))$ is a semilinear set.
@@ 4.1 Closure properties
We split the proof into several parts, each corresponding to the closure properties under one operation. Intersection. Let $L_{1}$ and $L_{2}$ be languages recognised respectively by $d_{1}$-dimensional $k_{1}$-ambiguous VASS $A_{1}$ and $d_{2}$-dimensional $k_{2}$-ambiguous VASS $A_{2}$. Language $L_{1}\cap L_{2}$ can be recognised by the standard synchronised product of $A_{1}$ and $A_{2}$. It is easy to observe that the product is a $(k_{1}\cdot k_{2})$-ambiguous VASS. Union. Let $L_{1}$ and $L_{2}$ be languages recognised by $d_{1}$-dimensional $k_{1}$-ambiguous VASS $A_{1}$ and $d_{2}$-dimensional $k_{2}$-ambiguous VASS $A_{2}$. The idea is to recognise $L_{1}\cup L_{2}$ by taking union VASS $A_{1}\cup A_{2}$, which is clearly $k_{1}+k_{2}$-ambiguous. Complementation. This comes from a general fact, that coverability languages are not closed under complementation. Language $L=a^{n}b^{\leq n}$ is recognised even by a deterministic VASS (hence also by one with bounded-ambiguity). On the other hand, in [CzerwinskiLMMKS18] it was shown, that every two disjoint coverability VASS [^1] In fact this result was shown for a wider class of Well-structured transition system (WSTS) languages $L_{1}$ and $L_{2}$ are regular separable, which means there exists a regular language $L_{3}$ such that $L_{1}\subseteq L_{3}$ and $L_{2}\cap L_{3}=\emptyset$. Because $L$ is a coverability VASS language its complement can be a coverability VASS language if and only if $L$ is a regular language. Clearly $L_{1}$ is not and this can be shown using the Pumping Lemma for regular languages.
@@
For each $k_{1},k_{2}\in\mathbb{N}_{+}$ there exists languages $L_{1},L_{2}$, which are respectively recognised by a $k_{1}$ and $k_{2}$ ambiguous VASS such that language $L_{1}\cup L_{2}$ is not recognised by a $n$-ambiguous VASS for $n\in[k_{1}+k_{2}-1]$.
@@
Let $k=k_{1}+k_{2}$. For $i\in[k]$ let us define language $U_{i}=\{a^{n_{1}}ba^{n_{2}}ba^{n_{3}}b\ldots a^{n_{k+1}}\mid n_{i}\geq n_{i+1}\}$. Observe, that $U_{i}$ is a language of an unambiguous VASS. Let $L_{1}=\bigcup_{i=1}^{k_{1}}U_{i}$ and $L_{2}=\bigcup_{i=k_{1}+1}^{k}$. By Lemma 4.1 languages $L_{1}$ and $L_{2}$ are respectively recognised by a $k_{1}$ and $k_{2}$ ambiguous VASS. By applying Lemma 3.1 to $k-1$ we get that language $L_{1}\cup L_{2}$ is not recognised by a $n$-ambiguous VASS for $n\in[k-1]$.
@@ 4.2 Expressiveness
Firstly, we show that, hierarchy of $k$-ambiguous VASS is strict.
@@
[…]
@@
Let $L=\{a^{n_{1}}ba^{n_{2}}ba^{n_{3}}b\ldots a^{n_{k+2}}\mid\exists_{1\leq i\leq k+1}n_{i}\geq n_{i+1}\}$. Because of Lemma 3.1 language $L$ is not recognised by a $k$-ambiguous VASS. On the other hand it is a union of $k+1$ unambiguous VASS languages. For $i\in[k+1]$ let us define $L_{i}=\{a^{n_{1}}ba^{n_{2}}ba^{n_{3}}b\ldots a^{n_{k+2}}\mid n_{i}\geq n_{i+1}\}$. Observe, that $L=\bigcup_{i=1}^{k+1}L_{i}$. Hence, by Lemma 4.1, $L$ is recognised by a $k+1$-ambiguous VASS.
@@
[…]
@@
[…]
$q_{1}$$q_{2}$$q_{3}$$q_{4}$$b$$a,+1$$b,-1$$a,b$$a,+1$$b,-1$
@@
On the other hand, assume, towards contradiction, that $L$ is recognised by a history-deterministic VASS. It is easy to see, that $L_{1}=a^{n>0}b^{\leq n}$ is a language of a deterministic VASS. Hence it is also history-deterministic. In [history-deterministic] it was shown, that history-deterministic languages are closed under union. Hence $L_{2}=L\cup L_{1}=\{a,b\}^{*}a^{n>0}b^{\leq n}$ is also history-deterministic. However, in [history-deterministic] it was also shown, that $L_{2}$ is not a history-deterministic language. Hence we get, that $L$ is not a history-deterministic language.
@@
[…]
$q_{1}$$q_{2}$$q_{3}$$b,(0,1,0)$$a,(0,-1,1)$$a,(0,0,0)$$a,(-1,0,0)$$a,(1,0,0)$$a,(0,2,-1)$
@@
Observe, that counters can nondeterministically decrease at each step. A run of M is a finite sequence $q_{1}(a_{1})\stackrel{{\scriptstyle t_{1}}}{{\longrightarrow}}q_{2}(a_{2})\stackrel{{\scriptstyle t_{2}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle t_{n-1}}}{{\longrightarrow}}q_{n}(a_{n})$. Given a configuration $q(u)$, the reachability set of $q(u)$ is the set of all configurations reachable from $q(u)$ via runs of M. We denote this set as $\textsc{Reach}(q(u))$. For simplicity we denote by $\textsc{Reach}(q)$ the set of configurations reachable from $q(\vec{0})$. It was shown in [lcm2] that the problem of deciding whether for a configuration $q(u)$ and LCM $M$ set $\textsc{Reach}(q(u))$ configuration is finite, is undecidable. Due to [DBLP:conf/concur/AlmagorY22] even if $u$ is always equal to $\vec{0}$ the problem is still undecidable. We call this problem 00-finite reach. We prove Theorem 1.1 by reducing from 00-finite reach. The proof is similar to the proofs of undecidability of regularity [regularity] and determinization [DBLP:conf/concur/AlmagorY22]. The rest of the section is devoted to the proof of Theorem 1.1.
Firstly, we present an overview of the proof. For each LCM $M_{1}$ with an initial state we create another LCM $M$ and initial state $l_{0}$ with the same answer to the 00-finite reach problem. Then, we define a language $L_{M,l_{0}}$, which intuitively encodes the correct runs of $M$. For technical reasons, namely because coverability VASSs are well-suited for recognizing languages similar to $a^{n}b^{\leq n}$ we encode the correct runs in reverse, that means from the final configuration to the initial one. Moreover, because we work with LCMs, it is better for us to work with the complement of $L_{M,l_{0}}$, that is $\overline{L_{M,l_{0}}}$. In such a way we get at the end a language for which Lemma 3.2 is useful. Then, the proof of Theorem 1.1 is split into three claims. Claim 15 states, that $\overline{L_{M,l_{0}}}$ is recognized by an effectively constructable $1$-dimensional VASS $A$. Claim 16 provides, that if Reach($l_{0}$) is finite than $L_{M,l_{0}}$ is a regular language. Finally, Claim 17 states, that if Reach($l_{0}$) is not finite then $L_{M,l_{0}}$ is not recognized by a VASS from the class of all boundedly-ambiguous VASS. All these claims give a direct reduction from 00-finite reach to deciding whether a language of a $1$-VASS belongs to class $C$ of languages, which contains all regular languages and is contained in the class of all boundedly-ambiguous VASS languages. Thus they conclude the proof of Theorem 1.1.
@@
Observe, that because each of the added transitions does not increase the sum of the counters, from $q_{1}$ we can go only to $q_{2}$ and later we can only stay in $q_{2}$ the answer for 00-finite reach is the same for both: $M_{1}$ and $M$. Hence, we proceed later with $M$. We encode each configuration $q(a_{1},a_{2},\ldots,a_{n})$ as a word over $\Sigma=Q\cup Z$ as $qz_{1}^{a_{1}}z_{2}^{a_{2}}\ldots z_{n}^{a_{n}}$. We use encodings of configurations to obtain an encoding of a run by concatenating encodings of its configurations. Finally, we define language $L_{M,l_{0}}=\{w^{r}\mid w\text{ is an encoding of a run from }l_{0}(0,0,\ldots,0)\}$.
@@
One can construct $1$-dimensional VASS recognizing $\overline{L_{M,l_{0}}}$.
@@
W.l.o.g. we assume, that there is at most one transition between each pair of states. We construct $A$ such that it accepts $w$ if and only if $w^{r}$ does not represent a valid run of $M$ from $l_{0}(0,0,\ldots,0)$. The idea is, that $A$ guesses the violation in the run represented by the word $w$. We have three types of violations. The first type is a control state violation, that is the run uses nonexisting transition or starts not in $l_{0}(0,0,\ldots,0)$. The second type is a counter violation, that we have invalid counter values between two consecutive configurations. The third violation is that we have invalid encoding of a configuration, that is we have two consecutive letters $z_{i}z_{j}$ such that $j>i$. To spot control violation for nonexisting transitions we will have gadget for each pair of states $p$ and $q$ such that there is no transition from $p$ to $q$ spotting infix of the form $z_{n}^{*}z_{n-1}^{*}\ldots z_{1}^{*}qz_{n}^{*}z_{n-1}^{*}\ldots z_{1}^{*}p$. Such gadget is just an NFA. Observe, that if a run start in $l_{0}(0,0,\ldots,0)$ then either the whole encoding of the run is equal to $l_{0}$ or suffix is of the form $pl_{0}$ for some $p\in Q$. Therefore to spot control violation, that the run does not start in $l_{0}(0,0,\ldots,0)$ we will have an NFA recognizing words such that suffix is not of the form $pl_{0}$ for some $p\in Q$ and the encoding is not equal to $l_{0}$. To spot counter violation we will have a 1-dimensional VASS for each transition and each counter spotting violation when firing this transition on this counter. Let us fix transition $t$ and counter $z_{i}$. Let transition $t$ be from state $p$ to state $q$. We have four possibilities for the operation performed by $t$ on the counter $z_{i}$. Therefore we need to spot infix of the form:
@@
    $z_{n}^{*}z_{n-1}^{*}\ldots z_{i}^{a}\ldots z_{1}^{*}qz_{n}^{*}z_{n-1}^{*}\ldots z_{i}^{b}\ldots z_{1}^{*}p$ where $a>b-1$ (equivalently $a\geq b$) if transition $t$ decrements counter $z_{i}$.
@@
All of the above can be done with $1$-dimensional VASS. To spot invalid encoding of a configuration for each $1\leq i<j\leq n$ we will have an NFA recognizing words with infix $z_{i}z_{j}$. As all the gadgets have one initial state in which we "ignore" some prefix of the word, we can join them and obtain one dimensional VASS with single initial configuration.
@@
If $\textsc{Reach}(l_{0})$ is infinite then $\overline{L_{M,l_{0}}}$ is not recognized by a boundedly-ambiguous VASS.
@@
As VASS $k$-ambiguous languages are closed under intersection with a regular language also $L_{1}=\overline{L_{M,l_{0}}}\cap(z_{1}^{*}q_{2})^{k+2}Z^{*}q_{1}(Q\cup Z\setminus q_{2})^{*}$ is recognized by a $k$-ambiguous VASS. The idea of this intersection is to get a language similar to the language presented in Lemma 3.2. Observe, that each $w\in L_{1}$ can be uniquely decomposed into $w=z_{1}^{n_{1}}q_{2}z_{1}^{n_{2}}q_{2}\ldots z_{1}^{n_{k+2}}q_{2}v$. We denote $v$ by $\textup{suff}(w)$. We define $L=\{\textup{suff}(w)\mid w\in L_{1}\}$. We define function $f$ on words from $L$ by setting $f(w)=0$ if $w^{r}$ encodes an incorrect run of $M$ from $l_{0}(0,0,\ldots)$ and otherwise $f(w)=n$ where $n$ is the maximal number such that $w=z_{1}^{n}v$ for some word $v$. Observe, that $L_{1}=\{z_{1}^{n_{1}}q_{2}z_{1}^{n_{2}}q_{2}z_{1}^{n_{3}}q_{2}\ldots z_{1}^{n_{k+2}}q_{2}w\mid w\in L,\exists_{1\leq i\leq k+1}n_{i}\geq n_{i+1}\lor n_{k+2}\geq f(w)\}$. This is because both transition from $q_{1}$ to $q_{2}$ and loop from $q_{2}$ to $q_{2}$ decrements the first counter and therefore for each $v\in L_{1}$ such that $v=z_{1}^{n_{1}}q_{2}z_{1}^{n_{2}}q_{2}z_{1}^{n_{3}}q_{2}\ldots z_{1}^{n_{k+2}}q_{2}w$ we have that $v^{r}$ encodes invalid run if and only if either $\textup{suff}(v)^{r}=w^{r}$ encodes an incorrect run (that is $f(w)=0\leq n_{k+2}$) or $v^{r}$ encodes a correct run but $f(w)\leq n_{k+2}$ (recall the definition of $f(w)$ in this case) or there exists $i\in[k+1]$ such that $n_{i}\geq n_{i+1}$. Moreover, notice that because $\textsc{Reach}(l_{0})$ is infinite and we have transitions moving values to the first counter from all the other counters in the state $q_{1}$ for each each $B\in\mathbb{N}$ there exist $B^{\prime}\in\mathbb{N}$ such that $B\leq B^{\prime}$ and a correct run $\rho_{B}$ of $M$ from $l_{0}(0,0,\ldots,0)$ to $q_{1}(B^{\prime},0,\ldots,0)$. Let $w_{B}$ be the encoding of this run. Observe, that $(z_{1}^{B^{\prime}}q_{2})^{k+2}w_{B}^{r}\in L_{1}$. Hence $w_{B}^{r}\in L$ and $f(w_{B}^{r})=B^{\prime}\geq B$. Hence $\sup(f)=\omega$ and hence, by Lemma 3.2, $L_{1}$ is not recognized by a $k$-ambiguous VASS. Contradiction.

### v2

@@ 3.5 Semilinear image
Now, we develop the second tool for showing that language is not recognised by a $k$-ambiguous VASS. Before formulating the tool we have to provide a few definitions. For any language $L\subseteq\{a,b\}^{*}$ such that for each $w\in L$ we have $_{b}(w)=l$ for some $l\in\mathbb{N}$ we define
@@
  • $L$ is recognised by $k$-ambiguous VASS $V$.
  • There exists $n\in\mathbb{N}$ such that for each $w\in L$ we have $_{b}(w)=n$.
@@
Proof 3.21.
The proof is based on Lemma 3.5 and the fact that set of solution of a system of linear Diophantine inequalities is semilinear. Notice first, that $L$ satisfies conditions of Lemma 3.5. Therefore we can apply Lemma 3.5 to $L$ and $n+1$ and decompose each accepting run in $V$ in the way presented in Lemma 3.5. Let us fix some $\alpha_{1},\ldots,\alpha_{n+1}$, $\alpha_{1}^{\prime},\ldots,\alpha_{n+1}^{\prime}$ and $\beta_{1},\ldots,\beta_{n+1}$, initial configuration $q(c)$ and accepting configuration $p(f)$. We call a run $\rho$ accepting with respect to $q(c)$ and $p(f)$ if $\rho$ is an accepting run of the VASS starting in $q(c)$ and ending in configuration $p(c^{\prime})$ such that $c^{\prime}\geq f$. Let $K$ be the following language:
@@
$|\beta_{n}|a_{n}+|\alpha_{n}|+|\alpha_{n}^{\prime}|-1,|\beta_{k+1}|a_{n+1}+|\alpha_{n+1}|+|\alpha_{n+1}^{\prime}|)\mid$
[…]
@@
Then, as shown in Lemma 4.1, $k$-ambiguous VASS languages are closed under intersection with regular languages. As mentioned below languages $K_{n}$ are regular for each $n\in\mathbb{N}$, the following corollary holds:
@@
Let $a,b\in\Sigma$, $L\subseteq\Sigma^{*}$ be a language recognised by a $k$-ambiguous VASS and for $n\in\mathbb{N}$ let $K_{n}\subseteq\{a,b\}^{*}$ be the language of words containing exactly $n$ letters $b$. Then for any $n\in N$ we have that $im(L\cap K_{n})$ is a semilinear set.
@@ 4.1 Closure properties
We split the proof into several parts, each corresponding to the closure properties under one operation. Intersection. Let $L_{1}$ and $L_{2}$ be languages recognised respectively by $d_{1}$-dimensional $k_{1}$-ambiguous VASS $A_{1}$ and $d_{2}$-dimensional $k_{2}$-ambiguous VASS $A_{2}$. Language $L_{1}\cap L_{2}$ can be recognised by the standard synchronised product of $A_{1}$ and $A_{2}$. It is easy to observe that the product is a $(k_{1}\cdot k_{2})$-ambiguous $(d_{1}+d_{2})$-VASS. Union. Let $L_{1}$ and $L_{2}$ be languages recognised by $d_{1}$-dimensional $k_{1}$-ambiguous VASS $A_{1}$ and $d_{2}$-dimensional $k_{2}$-ambiguous VASS $A_{2}$. The idea is to recognise $L_{1}\cup L_{2}$ by taking union VASS $A_{1}\cup A_{2}$, which is clearly $k_{1}+k_{2}$-ambiguous and $\max(d_{1},d_{2})$-dimensional.
Complementation. This comes from a general fact, that coverability languages are not closed under complementation. Language $L=a^{n}b^{\leq n}$ is recognised even by a deterministic VASS (hence also by one with bounded-ambiguity). On the other hand, in [CzerwinskiLMMKS18] it was shown, that every two disjoint coverability VASS [^1] In fact this result was shown for a wider class of Well-structured transition system (WSTS) languages $L_{1}$ and $L_{2}$ are regular separable, which means there exists a regular language $L_{3}$ such that $L_{1}\subseteq L_{3}$ and $L_{2}\cap L_{3}=\emptyset$. Because $L$ is a coverability VASS language its complement can be a coverability VASS language if and only if $L$ is a regular language. Clearly $L_{1}$ is not and this can be shown using the Pumping Lemma for regular languages.
@@
For each $k_{1},k_{2}\in\mathbb{N}_{+}$ there exist languages $L_{1},L_{2}$, which are respectively recognised by a $k_{1}$ and $k_{2}$ ambiguous VASS such that language $L_{1}\cup L_{2}$ is not recognised by an $n$-ambiguous VASS for $n\in[k_{1}+k_{2}-1]$.
@@
Let $k=k_{1}+k_{2}$. For $i\in[k]$ let us define language $U_{i}=\{a^{n_{1}}ba^{n_{2}}ba^{n_{3}}b\ldots a^{n_{k+1}}\mid n_{i}\geq n_{i+1}\}$. Observe, that $U_{i}$ is a language of an unambiguous VASS. Let $L_{1}=\bigcup_{i=1}^{k_{1}}U_{i}$ and $L_{2}=\bigcup_{i=k_{1}+1}^{k}$. By Lemma 4.1 languages $L_{1}$ and $L_{2}$ are respectively recognised by a $k_{1}$ and $k_{2}$ ambiguous VASS. By applying Lemma 3.1 to k1k-1 we get that language $L_{1}\cup L_{2}$ is not recognised by a $n$-ambiguous VASS for $n\in[k-1]$.
@@ 4.2 Expressiveness
Firstly, we show that the hierarchy of $k$-ambiguous VASS is strict.
@@
For every $k\in\mathbb{N}_{+}$ there exists language $L\in\textup{(k+1)-Amb}\setminus\textup{k-Amb}$.
@@
Let $L=\{a^{n_{1}}ba^{n_{2}}ba^{n_{3}}b\ldots a^{n_{k+2}}\mid\exists_{i\in[k+1]}\ n_{i}\geq n_{i+1}\}$. Because of Lemma 3.1 language $L$ is not recognised by a $k$-ambiguous VASS. On the other hand it is a union of $k+1$ unambiguous VASS languages. For $i\in[k+1]$ let us define $L_{i}=\{a^{n_{1}}ba^{n_{2}}ba^{n_{3}}b\ldots a^{n_{k+2}}\mid n_{i}\geq n_{i+1}\}$. Observe, that $L=\bigcup_{i=1}^{k+1}L_{i}$. Hence, by Lemma 4.1, $L$ is recognised by a $(k+1)$-ambiguous VASS.
@@
There exists language $L\in\textup{1-Amb}\setminus\textup{Det}$.
We show an even stronger Lemma 4.10. It is stronger because we have $\textup{Det}\subseteq\textup{Hist}$.
@@
There exists language $L\in\textup{1-Amb}\setminus\textup{Hist}$.
$q_{1}$$q_{2}$$q_{3}$$q_{4}$bba,+1a,+1b,1b,-1a,ba,ba,+1a,+1b,1b,-1
@@
On the other hand, assume, towards contradiction, that $L$ is recognised by a history-deterministic VASS. It is easy to see, that $L_{1}=a^{n>0}b^{\leq n}$ is a language of a deterministic VASS. Hence it is also history-deterministic. In [history-deterministic] it was shown, that history-deterministic VASS languages are closed under union. Hence $L_{2}=L\cup L_{1}=\{a,b\}^{*}a^{n>0}b^{\leq n}$ is also history-deterministic VASS language. However, in [history-deterministic] it was also shown, that $L_{2}$ is not a history-deterministic VASS language. Hence we get, that $L$ is not a history-deterministic VASS language.
@@
Observe, that because each word is read by only finite number of accepting runs, one can see bounded-ambiguity as an extension of the notion of determinism. A second extension of the determinism is history-determinism. In [history-deterministic] it was shown, that history-deterministic VASSs can express more than deterministic ones and less than nondeterministic ones. Up to now, there has been no comparison of the expressive power of history-deterministic and bounded-ambiguous VASSs. Now, we show, that these language classes are incomparable, which means that there exists a language recognised by an unambiguous VASS, which is not a language of history-deterministic VASS and language which is recognised by a history-deterministic VASS and it is not recognised by a $k$-ambiguous VASS for any $k\in\mathbb{N}_{+}$. Recall, that in Lemma 4.10 we have shown that there exists $L\in\textup{1-Amb}\setminus\textup{Hist}$. Hence it is enough to show the following Lemma:
$q_{1}$$q_{2}$$q_{3}$b,(0,1,0)b,(0,1,0)a,(0,1,1)a,(0,-1,1)a,(0,0,0)a,(0,0,0)a,(1,0,0)a,(-1,0,0)a,(1,0,0)a,(1,0,0)a,(0,2,1)a,(0,2,-1)
@@
Observe, that counters can nondeterministically decrease at each step. A run of M is a finite sequence $q_{1}(a_{1})\stackrel{{\scriptstyle t_{1}}}{{\longrightarrow}}q_{2}(a_{2})\stackrel{{\scriptstyle t_{2}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle t_{n-1}}}{{\longrightarrow}}q_{n}(a_{n})$. Given a configuration $q(u)$, the reachability set of $q(u)$ is the set of all configurations reachable from $q(u)$ via runs of M. We denote this set as $\textsc{Reach}(q(u))$. For simplicity we denote by $\textsc{Reach}(q)$ the set of configurations reachable from $q(\vec{0})$. It was shown in [lcm2] that the problem of deciding whether for a configuration $q(u)$ and LCM $M$ set $\textsc{Reach}(q(u))$ configuration is finite, is undecidable. Due to [DBLP:conf/concur/AlmagorY22] even if $u$ is always equal to $\vec{0}$ the problem is still undecidable. We call this problem 0-finite reach. We prove Theorem 1.1 by reducing from 0-finite reach. The proof is similar to the proofs of undecidability of regularity [regularity] and determinization [DBLP:conf/concur/AlmagorY22]. The rest of the section is devoted to the proof of Theorem 1.1.
Firstly, we present an overview of the proof. For each LCM $M_{1}$ with an initial state we create another LCM $M$ and initial state $l_{0}$ with the same answer to the 0-finite reach problem. Then, we define a language $L_{M,l_{0}}$, which intuitively encodes the correct runs of $M$. For technical reasons, namely because coverability VASSs are well-suited for recognising languages similar to $a^{n}b^{\leq n}$ we encode the correct runs in reverse, that means from the final configuration to the initial one. Moreover, because we work with LCMs, it is better for us to work with the complement of $L_{M,l_{0}}$, that is $\overline{L_{M,l_{0}}}$. In such a way we get at the end a language for which Lemma 3.2 is useful. Then, the proof of Theorem 1.1 is split into three claims. Claim 15 states, that $\overline{L_{M,l_{0}}}$ is recognised by an effectively constructible $1$-dimensional VASS $A$. Claim 16 provides, that if Reach($l_{0}$) is finite then $L_{M,l_{0}}$ is a regular language. Finally, Claim 17 states, that if Reach($l_{0}$) is not finite then $L_{M,l_{0}}$ is not recognised by a VASS from the class of all boundedly-ambiguous VASS. All these claims give a direct reduction from 0-finite reach to deciding whether a language of a $1$-VASS belongs to class $C$ of languages, which contains all regular languages and is contained in the class of all boundedly-ambiguous VASS languages. Thus they conclude the proof of Theorem 1.1.
@@
[…]
@@
One can construct $1$-dimensional VASS recognising $\overline{L_{M,l_{0}}}$.
@@
W.l.o.g. we assume, that there is at most one transition between each pair of states (otherwise we may split a state into several states, one for each incoming letter). We construct $A$ such that it accepts $w$ if and only if $w^{r}$ does not represent a valid run of $M$ from $l_{0}(0,0,\ldots,0)$. The idea is, that $A$ guesses the violation in the run represented by the word $w$. We have three types of violations. The first type is a control state violation, that is the run uses nonexisting transition or does not start in $l_{0}(0,0,\ldots,0)$. The second type is a counter violation, that we have invalid counter values between two consecutive configurations. The third violation is that we have invalid encoding of a configuration, that is we have two consecutive letters $z_{i}z_{j}$ such that $j>i$. To spot control violation for nonexisting transitions we will have gadget for each pair of states $p$ and $q$ such that there is no transition from $p$ to $q$ spotting infix of the form $z_{n}^{*}z_{n-1}^{*}\ldots z_{1}^{*}qz_{n}^{*}z_{n-1}^{*}\ldots z_{1}^{*}p$. Such gadget is just an NFA. Observe, that if a run start in $l_{0}(0,0,\ldots,0)$ then either the whole encoding of the run is equal to $l_{0}$ or suffix is of the form $pl_{0}$ for some $p\in Q$. Therefore to spot control violation, that the run does not start in $l_{0}(0,0,\ldots,0)$ we will have an NFA recognising words such that suffix is not of the form $pl_{0}$ for some $p\in Q$ and the encoding is not equal to $l_{0}$. To spot counter violation we will have a 1-dimensional VASS for each transition and each counter spotting violation when firing this transition on this counter. Let us fix transition $t$ and counter $z_{i}$. Let transition $t$ be from state $p$ to state $q$. We have four possibilities for the operation performed by $t$ on the counter $z_{i}$. Therefore we need to spot infix of the form:
@@
    $z_{n}^{*}z_{n-1}^{*}\ldots z_{i}^{a}\ldots z_{1}^{*}qz_{n}^{*}z_{n-1}^{*}\ldots z_{i}^{b}\ldots z_{1}^{*}p$ where a>b1a>b-1 (equivalently $a\geq b$) if transition $t$ decrements counter $z_{i}$.
@@
All of the above can be done with $1$-dimensional VASS. To spot invalid encoding of a configuration for each $1\leq i<j\leq n$ we will have an NFA recognising words with infix $z_{i}z_{j}$. As all the gadgets have one initial state in which we "ignore" some prefix of the word, we can join them and obtain one dimensional VASS with single initial configuration.
@@
If $\textsc{Reach}(l_{0})$ is infinite then $\overline{L_{M,l_{0}}}$ is not recognised by a boundedly-ambiguous VASS.
@@
As $k$-ambiguous VASS languages are closed under intersection with a regular language also $L_{1}=\overline{L_{M,l_{0}}}\cap(z_{1}^{*}q_{2})^{k+2}Z^{*}q_{1}(Q\cup Z\setminus q_{2})^{*}$ is recognised by a $k$-ambiguous VASS. The idea of this intersection is to get a language similar to the language presented in Lemma 3.2. Observe, that each $w\in L_{1}$ can be uniquely decomposed into $w=z_{1}^{n_{1}}q_{2}z_{1}^{n_{2}}q_{2}\ldots z_{1}^{n_{k+2}}q_{2}v$. We denote $v$ by $\textup{suff}(w)$. We define $L=\{\textup{suff}(w)\mid w\in L_{1}\}$. We define function $f$ on words from $L$ by setting $f(w)=0$ if $w^{r}$ encodes an incorrect run of $M$ from $l_{0}(0,0,\ldots,0)$ and otherwise $f(w)=n$ where $n$ is the maximal number such that $w=z_{1}^{n}v$ for some word $v$. Observe, that $L_{1}=\{z_{1}^{n_{1}}q_{2}z_{1}^{n_{2}}q_{2}z_{1}^{n_{3}}q_{2}\ldots z_{1}^{n_{k+2}}q_{2}w\mid w\in L,\exists_{1\leq i\leq k+1}n_{i}\geq n_{i+1}\lor n_{k+2}\geq f(w)\}$. This is because both transition from $q_{1}$ to $q_{2}$ and loop from $q_{2}$ to $q_{2}$ decrements the first counter and therefore for each $v\in L_{1}$ such that $v=z_{1}^{n_{1}}q_{2}z_{1}^{n_{2}}q_{2}z_{1}^{n_{3}}q_{2}\ldots z_{1}^{n_{k+2}}q_{2}w$ we have that $v^{r}$ encodes invalid run if and only if either $\textup{suff}(v)^{r}=w^{r}$ encodes an incorrect run (that is $f(w)=0\leq n_{k+2}$) or $v^{r}$ encodes a correct run but $f(w)\leq n_{k+2}$ (recall the definition of $f(w)$ in this case) or there exists $i\in[k+1]$ such that $n_{i}\geq n_{i+1}$. Moreover, notice that because $\textsc{Reach}(l_{0})$ is infinite and we have transitions moving values to the first counter from all the other counters in the state $q_{1}$ for each each $B\in\mathbb{N}$ there exist $B^{\prime}\in\mathbb{N}$ such that $B\leq B^{\prime}$ and a correct run $\rho_{B}$ of $M$ from $l_{0}(0,0,\ldots,0)$ to $q_{1}(B^{\prime},0,\ldots,0)$. Let $w_{B}$ be the encoding of this run. Observe, that $(z_{1}^{B^{\prime}}q_{2})^{k+2}w_{B}^{r}\in L_{1}$. Hence $w_{B}^{r}\in L$ and $f(w_{B}^{r})=B^{\prime}\geq B$. Hence $\sup(f)=\omega$ and hence, by Lemma 3.2, $L_{1}$ is not recognised by a $k$-ambiguous VASS. Contradiction.